your plan to finance a toyota camry that sells for $32,000 with a loan that carries a 3% APR for 5 years (60 months). what is your monthly payment?
Answer #1

Monthly payment = ( interest rate / 12 months) *(principal amount ) / (1 -(1+ interest rate /12)(-number of months))

Interest rate = 0.03

principal amount =$ 32000

number of months = 60

year = 12

monthly payment = (0.03 /12) * (32000) / 1- (1+ 0.03 / 12) (-60)

= 32000 * 0.0025 / 1-(1.0025)(-60)

= 31999.99 / 1 -0.860869

= 80 / 0.1391

= 575.12